Two concentric conducting spheres of radii a = 5 cm and b = 10 cm have a potential difference V =
100 V between them. What is the charge carried by each sphere?
Answer #1
The potential difference between them
V = keQ (b-a)/ (ab)
==> 100 = 9*10^9*Q*5*10^-2/50*10^-4
==> Q = 111.1*10^-11 C
Q = 1.1nC
Ans: 1.1 nC
Answer #2
let be q

b-a/ab * kq = 100

kq = 1000

q = 111.1 * 10^(-9)

111.1 nC